Grade 11Algebra

if a and b are integers such that alog2502 + blog250b = 3 Find a+2b?

Rogel Luyun
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
log250 (2^a * b^b) = 3
 
2^a * b^b = 250 ^3
 
2^a *b^b = 2 * 5 * 5 * 5 * 2 * 5 * 5 * 5 * 2 * 5 * 5 * 5
 
Hope it helps
 
As it is some thing missing in mysolution